Find f(3), given that f is a linear function with f(1)=2 and f(4)=-7
Answer by rothauserc(4718)

m = (-7 -2) / (4 - 1) = -3
2 = -3 + b
b = 5
f(3) = -3*3 + 5 = -4
note that the linear function is f(x) = -3x + 5